# Ivan Gatica and Alfonso Bustamante

### v2

@@ 1 $\sigma$-Sets and $\sigma$-Antisets
We must observe that given the definition of the fusion operator $\oplus$ in [6] it is clear that it is commutative and therefore if $B$ is an $\sigma$-antiset of $A$, then it will be necessary that $A$ is also the $\sigma$-antiset of $B$. On the other hand, following Blizard notation, [4] p. 347, we will denote $B$ the $\sigma$-antiset of $A$ as $B=A^{-}$, in this way we will have that $A=(A^{-})^{-}$.
@@
where both chains have the linear $\in$-root property and are totally different, i.e. they do not have a link-intersection.
@@
Let $A$ and $B$ be $\sigma$-sets. The Generated space by $A$ and $B$ is given by
@@
Let us recall a few things about the fusion operator $\oplus$. To begin this brief analysis, we must observe that given $x$, $y$ two $\sigma$-sets, if $\{x\}\cup\{y\}=\emptyset$ then it will be said that $y$ is the antielement of $x$ and $x$ the antielement of $y$, where the union of pairs $\cup$ axiomatized within the theory of $\sigma$-sets is used, in particular in the completion axioms A and B, which we will call annihilation axioms from now on.
@@
Let $x$ be an element of some $\sigma$-set, then we will denote as $x^{\ast}$ the antielement of $x$, in the case that it exists.
Now we move on to define the new operations with $\sigma$-sets which will help us define the fusion of $\sigma$-sets.
@@
Let $A$ and $B$ be $\sigma$-sets, then we define the $\ast$-intersection of $A$ with $B$ by
@@
Let $A=\{1,2,3^{\ast},4\}$ and $B=\{2,3,4^{\ast}\}$, then we have that:
@@
Let $A$ be $\sigma$-set, then $A\widehat{\cap}A=\emptyset$.
@@
Regarding Theorem 1.6, we can observe that given a $\sigma$-set $A$ the $\sigma$-set theory does not allow the coexistence of a $\sigma$-element $x$ and its $\sigma$-antielement in the same $\sigma$-set $A$, and this is because $A$ is a $\sigma$-set. However, since $\sigma$-set theory is a $\sigma$-class theory, one can find the $\sigma$-elements together with the $\sigma$-antielements coexisting without problems in what we call the proper $\sigma$-class, in this way one will have that $\{x,x^{\ast}\}$ is a proper $\sigma$-class and not a $\sigma$-set.
@@
Let $A$ be $\sigma$-set, then $A\widehat{\cap}\emptyset=\emptyset$ and ^$\emptyset\widehat{\cap}A=\emptyset$.
@@
Now suppose that $A\widehat{\cap}\emptyset\neq\emptyset$, then there exists an $x\in A$ such that $x^{\ast}\in\emptyset$, which is a contradiction, hence $A\widehat{\cap}\emptyset=\emptyset$. On the other hand, ^$\emptyset\widehat{\cap}A\subseteq\emptyset$ thus we will have that ^$\emptyset\widehat{\cap}A=\emptyset$. ∎
@@
Let $A$ and $B$ be $\sigma$-sets, then we define the $\ast$-difference between $A$ y $B$ by
@@
Let $A$ be $\sigma$-set. Then $A\divideontimes A=A$.
@@
Let $A$ be $\sigma$-set. Then $A\divideontimes\emptyset=A$ and $\emptyset\divideontimes A=\emptyset$.
@@
Let $A$ and $B$ be $\sigma$-sets, then we define the fusion of $A$ and $B$ by
@@
It is clear that the fusion of $\sigma$-sets is commutative by definition. Now let us show an example
@@
Let $A$ be $\sigma$-set, then $A\oplus\emptyset=\emptyset\oplus A=A$.
@@
Second we will show that $\emptyset\oplus A=A$. By definition we will have that,
@@
As we said before the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [6, 3, 2] this operation is not associative.
@@ 2 Generated space
It is important to note that from the perspective of $\sigma$-sets we have that $\emptyset=\emptyset^{-}=\emptyset^{i}_{j}$ with $i\in\{0,1,2,3,4,5,6,7\}$ and $j\in\{0,1,2,3\}$, where the difference of the $\sigma$-emptysets $\emptyset^{i}_{j}$ is given by annihilation, which come from the equation $A\oplus A^{-}=\emptyset$.
@@
Finally as a theoretical result we have a cardinal theorem:
@@
We should also note that we have obtained other cardinal results for the integer space $3^{A}$ with $\left|A\right|\in\{0,1,2,3,4,5\}$. The cardinal results are the following:
@@
Let $A$ be $\sigma$-set such that $|A|=n$, then $\left|3^{A}\right|=\left|\left\langle 2^{A},2^{A^{-}}\right\rangle\right|=3^{n}$.
On the other hand, as we have already said, we are going to change the notation of $1_{\Theta}$ to $1_{0}$, in this way we will have the set 00-natural numbers defined as follows:
@@
In this case the generated space becomes a meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ which can be ordered graphically as shown in figure 2.
@@
Now if we count the number of elements that the meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ has, we will have that they are 36, where the prime decomposition of this number is $36=2^{2}\cdot 3^{2}$ which is equivalent to the following multiplication of cardinals $36=2^{|A|}\cdot 3^{|B|}$, from here we can obtain the following conjecture:
@@ 3 Algebraic structure of integer space $3^{A}$
From here it is clearly seen that the conditions $(1)$, $(2)$ and $(3)$ of theorem 3.1 are satisfied, where the condition $(4)$ is obvious by definition.
We must clarify that since $\sigma$-set $\emptyset=\emptyset^{-}$, and also $\emptyset=\emptyset^{0}_{0}=\emptyset^{1}_{1}=\emptyset^{2}_{1}=\emptyset^{3}_{2}$, from here we have the condition $(2)$ and the difference is in another dimension, the dimension of annihilation. Here we must clarify that the fusion operation $\oplus$ is not associative. Let $X=\{1^{\ast},2^{\ast}\}$, $Y=\{1,2\}$ and $Z=\{1\}$ then we will have that ($(\{1^{\ast},2^{\ast}\}\oplus\{1,2\})\oplus\{1\}=\emptyset\oplus\{1\}=\{1\}$
@@ 4 $\sigma$-Sets Equations
It is clear that if a $\sigma$-set $A$ is integer, then by definition there exists the integer space $3^{A}$. We should also note that if $A$ is an integer $\sigma$-set, then $\left[A\cup A^{-}\right]$ is a proper $\sigma$-class, for example, consider $A=\{1,2\}$, then $\left[A\cup A^{-}\right]=\left[1,2,1^{\ast},2^{\ast}\right]$, is a proper $\sigma$-class.
@@
Let $A$ be an integer $\sigma$-set such that $|A|=n$, then $X$ is said to be a $\sigma$-set variable of $3^{A}$, if and only if
@@
We must remember that the structure in general is not associative, therefore we cannot freely use this property, so to find the solution to the equation we must develop a previous theorem. To develop this theorem we will assume that for every integer $\sigma$-set $A$ then the generated space $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also that $3^{A}$ satisfies the conjecture 3.2.
@@
Let $A$ be an entire $\sigma$-set, $X$ be a $\sigma$-set variable of $3^{A}$ and $M\in 3^{A}$, then by lemma 4.5 we have that $X\oplus M=X\cup M$, with $X\cap M=\emptyset$.
@@
Now that theorem 4.7 has been proved, we can solve some $\sigma$-set equation for the integer $\sigma$-set $A=\{1,2\}$, since the generated space is effectively equal to $3^{A}$, that is, $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also $3^{A}$ is a non-associative abelian loop.
Let $A=\{1,2\}$ be an integer set and $M,N\in 3^{A}$, then the equation
@@
Let $A=\{1,2\}$ be an integer $\sigma$-set, $M=\{1,2^{\ast}\}$ and $N=\{1\}$, then the equation of a $\sigma$-set variable
@@
Before we conclude our study of $\sigma$-sets and $\sigma$-antisets we are going to present a conjecture regarding the possible solutions of a particular type of equations of a $\sigma$-set variable.
Conjecture 4.9.
Let $A$ be an integer $\sigma$-set, $X$ a $\sigma$-set variable of $3^{A}$, and $M,N\in 3^{A}$, then a solution of the equation
$X\oplus M=N,$
is $S=N\oplus M^{-}$.
Let us look into the following case, where $M=\{1^{*}\}$ and $N=\{1\}$, so for the equation
$X\oplus\{1^{*}\}=\{1\},$
there is no element $x$ such that $\{x,1^{*}\}=\{1\}$. This lead us into a more general case where $N\subseteq M^{-}$ or $N\supset M^{-}$, being the conjecture false, which take us into establishing the following condition in order for this type of equation being solvable.
@@
With this, let us conclude with a bounded theorem to find some solutions of the $\sigma$-set equation.

### v3

@@ 1 $\sigma$-Sets and $\sigma$-Antisets
We must observe that given the definition of the fusion operator $\oplus$ in [5] it is clear that it is commutative and therefore if $B$ is an $\sigma$-antiset of $A$, then it will be necessary that $A$ is also the $\sigma$-antiset of $B$. On the other hand, following the Blizard notation, [3] p. 347, we will denote $B$ the $\sigma$-antiset of $A$ as $B=A^{-}$, in this way we will have $A=(A^{-})^{-}$.
@@
where both $\epsilon$-chains have the linear $\epsilon$-root property and are totally different, i.e. they do not have a link-intersection. These definitions can be found in [5] Definition 3.13, 3.14 and 3.16.
@@
Let $A$ and $B$ be two $\sigma$-sets. The Generated space by $A$ and $B$ is given by
@@
Let us recall a few things about the fusion operator $\oplus$. In this brief analysis, we must observe that given $x$, $y$ two $\sigma$-sets, if $\{x\}\cup\{y\}=\emptyset$ then it will be said that $y$ is the antielement of $x$ and $x$ the antielement of $y$, where the union of pairs $\cup$ axiomatized within the theory of $\sigma$-sets is used, in particular in the completion axioms A and B, which we will call annihilation axioms from now on.
@@
Let $x$ be an element of some $\sigma$-set, then we will denote by $x^{\ast}$ the anti-element of $x$, if it exists.
Now we move on to define the new operations with $\sigma$-sets which will help us define the fusion of $\sigma$-sets $\oplus$.
@@
Let $A$ and $B$ be two $\sigma$-sets, then we define the $\ast$-intersection of $A$ with $B$ by
@@
Let $A=\{1,2,3^{\ast},4\}$ and $B=\{2,3,4^{\ast}\}$ be two $\sigma$-sets, then we have that:
@@
Let $A$ be a $\sigma$-set, then $A\widehat{\cap}A=\emptyset$.
@@
Example 1.7.
Let $A=\{1,2,3^{*},4\}$, then
$A\widehat{\cap}A=\{1,2,3^{*},4\}\widehat{\cap}\{1,2,3^{*},4\},$
$A\widehat{\cap}A=\{x\in\{1,2,3^{*},4\}:x^{*}\in\{1,2,3^{*},4\}\},$
$A\widehat{\cap}A=\emptyset.$
Regarding Theorem 1.6, we can observe that given a $\sigma$-set $A$, the $\sigma$-set theory does not allow the coexistence of a $\sigma$-element $x$ and its $\sigma$-antielement in the same $\sigma$-set $A$, and this is because $A$ is a $\sigma$-set. However, since $\sigma$-set theory is a $\sigma$-class theory, one can find the $\sigma$-elements together with the $\sigma$-antielements coexisting without problems in what we call the proper $\sigma$-class, in this way one will have that $\{x,x^{\ast}\}$ is a proper $\sigma$-class and not a $\sigma$-set.
@@
Let $A$ be a $\sigma$-set, then $A\widehat{\cap}\emptyset=\emptyset$ and ^$\emptyset\widehat{\cap}A=\emptyset$.
@@
Now suppose that $A\widehat{\cap}\emptyset\neq\emptyset$, then there exists an $x\in A$ such that $x^{\ast}\in\emptyset$, which is a contradiction, hence $A\widehat{\cap}\emptyset=\emptyset$. On the other hand, ^$\emptyset\widehat{\cap}A\subseteq\emptyset$ thus we will have to ^$\emptyset\widehat{\cap}A=\emptyset$. ∎
@@
Let $A$ and $B$ be two $\sigma$-sets, then we define the $\ast$-difference between $A$ y $B$ by
@@
Let $A$ be a $\sigma$-set. Then $A\divideontimes A=A$.
@@
Let $A$ be a $\sigma$-set. Then $A\divideontimes\emptyset=A$ and $\emptyset\divideontimes A=\emptyset$.
@@
Let $A$ and $B$ be two $\sigma$-sets, then we define the fusion of $A$ and $B$ by
@@
It is clear that the fusion of $\sigma$-sets is commutative by definition. Now, let us show an example
@@
Let $A$ be a $\sigma$-set, then $A\oplus\emptyset=\emptyset\oplus A=A$.
@@
Second, we will show that $\emptyset\oplus A=A$. By definition we will have that,
@@
Example 1.18.
Let $X=\{1,2,3\}$, $A=\{1,2\}$ and $B=\{2,3\}$, it is clear that $A,B\in 2^{X}$. Now we apply the fusion operator $\oplus$.
$A\oplus B=\{x:x\in A\divideontimes B\vee x\in B\divideontimes A\},$
$A\oplus B=\{x:x\in A\vee x\in B\},$
$A\oplus B=A\cup B=\{1,2,3\}.$
@@
Example 1.20.
Let $X=\{1,2,3,4\}$ and $A=\{1,2,3\}$, it is clear that $A\in 2^{X}$. Now we apply the fusion operator $\oplus$.
$A\oplus X=\{x:x\in A\divideontimes X\vee x\in X\divideontimes A\},$
$A\oplus B=\{x:x\in A\vee x\in X\},$
$A\oplus X=A\cup X=\{1,2,3,4\}=X.$
As we said before, the fusion of $\sigma$-sets $\oplus$ is commutative by definition but as we demonstrated in [5, 2, 1] this operation is not associative.
@@ 2 Generated space
It is important to note that from the perspective of $\sigma$-sets we have that $\emptyset=\emptyset^{-}=\emptyset^{i}_{j}$ with $i\in\{0,1,2,3,4,5,6,7\}$ and $j\in\{0,1,2,3\}$, where the difference of the $\sigma$-emptysets $\emptyset^{i}_{j}$ is given by annihilation, which comes from equation $A\oplus A^{-}=\emptyset$.
@@
Finally, as a theoretical result, we have a cardinal theorem:
@@
We should also note that we have obtained other cardinal results for the integer space $3^{A}$ with $\left|A\right|\in\{0,1,2,3,4,5\}$. The cardinal results are as follows:
@@
Let $A$ be a $\sigma$-set such that $|A|=n$, then $\left|3^{A}\right|=\left|\left\langle 2^{A},2^{A^{-}}\right\rangle\right|=3^{n}$.
On the other hand, as we have already said, we are going to change the notation of $1_{\Theta}$ to $1_{0}$, in this way we will have the $\sigma$-set of 00-natural numbers defined as follows:
@@
In this case, the generated space becomes a meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ which can be ordered graphically as shown in figure 2.
@@
Now, if we count the number of elements that the meta-space generated by $A=\{1_{0},2_{0}\}$ and $B=\{1,2\}$ has, we will find that they are 36, where the prime decomposition of this number is $36=2^{2}\cdot 3^{2}$ which is equivalent to the following multiplication of cardinals $36=2^{|A|}\cdot 3^{|B|}$, from where we can obtain the following conjecture:
@@ 3 Algebraic structure of integer space $3^{A}$
From here it is clearly seen that conditions $(1)$, $(2)$, and $(3)$ of theorem 3.1 are satisfied, where the condition $(4)$ is obvious by definition.
We must clarify that since $\sigma$-set $\emptyset=\emptyset^{-}$, and also $\emptyset=\emptyset^{0}_{0}=\emptyset^{1}_{1}=\emptyset^{2}_{1}=\emptyset^{3}_{2}$, from here we have condition $(2)$ and the difference is in another dimension, the dimension of annihilation. Here we must clarify that the fusion operation $\oplus$ is not associative. Let $X=\{1^{\ast},2^{\ast}\}$, $Y=\{1,2\}$ and $Z=\{1\}$ then we will have that ($(\{1^{\ast},2^{\ast}\}\oplus\{1,2\})\oplus\{1\}=\emptyset\oplus\{1\}=\{1\}$
@@ 4 $\sigma$-Sets Equations
It is clear that if a $\sigma$-set $A$ is integer, then by definition there exists the integer space $3^{A}$. We should also note that if $A$ is an integer $\sigma$-set, then $\left[A\cup A^{-}\right]$ is a proper $\sigma$-class, for example, consider $A=\{1,2\}$, then $\left[A\cup A^{-}\right]=\left[1,2,1^{\ast},2^{\ast}\right]$, is a proper $\sigma$-class. We must observe that $\sigma$-set theory [5] is a theory of $\sigma$-classes, where $\sigma$-sets are characterized by axioms. We must also note that a proper $\sigma$-class is a $\sigma$-class that is not a $\sigma$-set. This difference is essential to give rise to the existence of antielements along with their respective elements.
@@
Let $A$ be a integer $\sigma$-set such that $|A|=n$, then $X$ is said to be a $\sigma$-set variable of $3^{A}$, if and only if
@@
We must remember that the structure in general is not associative, therefore we cannot freely use this property, so to find the solution to the equation we must develop a previous theorem. To develop this theorem we will assume that for every integer $\sigma$-set $A$ the generated space is $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also that $3^{A}$ satisfies conjecture 3.2.
@@
Let $A$ be an integer $\sigma$-set, $X$ be a $\sigma$-set variable of $3^{A}$ and $M\in 3^{A}$, then by lemma 4.5 we have that $X\oplus M=X\cup M$, with $X\cap M=\emptyset$.
@@
Now, after theorem 4.7 has been proved, we can solve some $\sigma$-set equation for the integer $\sigma$-set $A=\{1,2\}$, since the generated space is effectively equal to $3^{A}$, that is, $\left\langle 2^{A},2^{A^{-}}\right\rangle=3^{A}$, and also $3^{A}$ is a non-associative abelian loop.
Let $A=\{1,2\}$ be an integer set and $M,N\in 3^{A}$, with $M\widehat{\cap}N=\emptyset$, then the equation
@@
Let $A=\{1,2\}$ be an integer $\sigma$-set, $M=\{1,2^{\ast}\}$ and $N=\{1\}$, with $M\widehat{\cap}N=\emptyset$, then the equation of a $\sigma$-set variable
@@
Note that if $M\widehat{\cap}N=\emptyset$ then the $\sigma$-set equation has a solution, but otherwise the $\sigma$-set equation has an empty solution.
Example 4.9.
Let $A=\{1,2\}$ be an integer $\sigma$-set, $M=\{1^{\ast}\}$ and $N=\{1\}$, with $M\widehat{\cap}N=\{1^{*}\}$, then the equation of a $\sigma$-set variable
$X\oplus\{1^{\ast}\}=\{1\}$
There is no solution.
$X\oplus\{1^{\ast}\}=\{1\}\setminus\oplus\{1\},$
$(X\oplus\{1^{\ast}\})\oplus\{1\}=\{1\}\oplus\{1\},$
$X=\{1\},$
which is a contradiction, because
$\{1\}\oplus\{1^{\ast}\}=\{1\},$
=$\emptyset=\{1\}.$
@@
With this in mind, let us conclude with a bounded theorem to find some solutions of the $\sigma$-set equation.
@@
5 Conclusions
One of the first conclusions we can draw is that the fusion operator $\oplus$ for $\sigma$sets is equivalent to the union operator for sets within the context of the set of parts $2^{A}$, which allows us to deduce that the fusion of $\sigma$-sets is an extension of the union for the generated space.
The fact that the integer space $3^{A}$ presents a cardinal of power 3, is very important for the development of the theory of transfinite numbers, since in general the power set $2^{A}$ that goes to the power of 2 is used; in this way our results can serve as an impetus for the development of the theory of transfinite numbers.
We can also conclude that the algebraic structure of the integer space $3^{1,2}$ is a loop, which leads us to conjecture that the integer space in general has a loop structure. This fact is relevant to $\sigma$-set theory since, if it were so, it would show that the fusion operator $\oplus$ is not associative which is relevant for solving set equations.
As a final conclusion, we can state that we can generate $\sigma$-set equations given the existence of inverses for the fusion operator $\oplus$ in the integer space, but in the general case, solutions are not given, so a condition must be imposed on the $\sigma$-sets of the equation. We have not yet conducted a detailed study on the number of solutions to each set equation, leaving this study for future research.
To see more works in which antisets or $\sigma$-antiset are used or in which equation $A\cup B=\emptyset$ is described, visit the references [1], [2], [4], [5], [6].